# DrawTalking: Building Interactive Worlds by Sketching and Speaking

### v3

@@ Abstract.
We introduce DrawTalking, an approach to building and controlling interactive worlds by sketching and speaking. It emphasizes user control and flexibility, and gives programming-like capability without requiring code. We built a prototype to demonstrate it.
An early open-ended study shows the mechanics resonate and are applicable to many creative-exploratory use cases, with the potential to inspire and inform research in future natural interfaces for creative exploration and authoring.
@@
ccs: Human-centered computingccs: Human-centered computing Graphical user interfacesccs: Human-centered computing Natural language interfacesccs: Human-centered computing Interaction techniquesccs: Human-centered computing Interactive systems and tools
@@
[Controls by sketching and speaking]This figure showcases the DrawTalking sketching + speaking controls. (1) On the left is a real-world photo of the user labeling a sketch on the digital canvas. The canvas contains a few user-drawn rectangle objects. The user narrates ”and I’m also going to have a rectangle,” which shows-up in a text transcript. Simultaneously, the user holds one of the rectangles with their finger and tapping the word ”rectangle” in the aforementioned transcript. This labels the user’s sketch with the word ”rectangle.” This is important because now the system knows what the name user’s object is, meaning it can now be controlled with more speech commands. (2) On the right is a mock-up of an example scenario; 3 panels. Panel 1: 3 lily pads have been drawn, as well as a frog. A caption reads a speech command: ”Forever this frog hops to a lily.” The user’s finger points to the frog to indicate which frog is being commanded. Dotted arrows indicate that the frog will now hop between the lilies once the command stops. Panel 2: Now a butterfly and some flowers have been drawn. The new command caption reads, ”The butterfly follows the frog.” Panel 3: The command caption reads ”This butterfly is slow” and the user’s finger is touching the butterfly, meaning the butterfly’s speed will now be decreased. Dotted arrows are now indicating that the butterfly has been moving around, following the frog as previously commanded. Slowing the butterfly does not require redoing the previous command.
@@ 1. Introduction
Sketching while speaking aids innovation, thinking, and communication — with applications in animation, game design, education, engineering, rapid prototyping, storytelling, and many other creative and spontaneous activities (Fan et al., 2023; Tversky, 2011). The combination enables us to think about and share anything through make-believe – including things that do not or cannot exist. We achieve this by assigning representations (sketches) to semantic concepts (objects, behaviors, properties) (Turner, 2016). For example we might suspend our disbelief so a square represents anything from a house, to a playing card, a map, dog, or person.
@@
This work is an attempt to realize a style of spontaneous interaction that seamlessly integrates sketching and talking-out-loud to build interactive, explorable worlds and simulations, thus greatly increasing our range of computational expression (Victor, 2014). Prior works in interactive sketching (Sutherland, 1963; Kazi et al., 2014a; Suzuki et al., 2020; Perlin et al., 2018; Saquib et al., 2021), language/AI-mediated interfaces (Winograd, 1972; Bolt, 1980; Liu et al., 2023; Xia et al., 2023; Huang et al., 2020; Xia, 2020; Subramonyam et al., 2018; Laput et al., 2013), and visual programming-adjacent interfaces or games (Sutherland, 1966; Resnick et al., 2009; Landay, 1996; Perlin and Goldberg, 1996; cycling74, 2018; Little Big Planet 2, 2011; Dreams, 2020) has laid valuable groundwork. However, they often require content to be pre-built or assume that users have predetermined goals, limiting the user’s potential to be spontaneous. Tools focus on generating a specific output rather than facilitating an ongoing creative process (Compton and Mateas, 2015). They might enforce given representations of objects (e.g. realistic or specific sketch recognition). They might feature complex UI, and in the case of programming-oriented tools, require explicit programming knowledge (whether via text, nodes, or blocks).
@@
In our prototype system, DrawTalking, users speak while freehand-sketching to create, control, and iterate on interactive visual mechanisms, simulations, and animations. Via speech and direct manipulation, the user names their sketches to provide semantic information to the system, and narrates desired behaviors and rules to influence the interactive simulation, as if by explaining to others or telling a story. This is inspired by the way in which people might speak and/or write text to describe objects as they are being drawn. The key insight is that in the context of narration, the machine can use this information directly without guessing-at the user’s intended meaning behind their drawings. As a result, users can build and control interactive worlds by simple combinations of sketching and speaking. DrawTalking requires no preparation or staging, and supports the user in making changes to content and behavior anytime. By design, we balance AI-automation with user-direction such that the user is the one who chooses how to represent content, when to do operations, and what logic to define.
@@
  • DrawTalking: a novel sketching+speech interaction that balances user-agency with machine automation.
@@
  • An early qualitative study of DrawTalking that reveals its use for emergent creativity and playful ideation
@@ 2.1. Natural Language-Adjacent Interfaces
Systems such as SHRDLU (Winograd, 1972) and Put That There (Bolt, 1980) pioneered the vision of employing natural language to communicate with computers. Due to recent advances in speech recognition and natural language understanding, the popularity of this interaction modality has exploded, and has been used in a wide range of domains. For example, VoiceCut (Kim et al., 2019) and PixelTone (Laput et al., 2013) allow users to speak short phrases or sentences to perform desired operations in image editing applications, but these applications are heavily domain-specific. Tools like WordsEye (Coyne and Sproat, 2001), Scones (Huang et al., 2020), and CrossPower (Xia, 2020) enable scene generation or content editing via language, and interfaces such as Visual Captions (Liu et al., 2023), RealityTalk (Liao et al., 2022), and CrossTalk (Xia et al., 2023) use language structure to make content appear during talks or conversations. However, many of these interfaces tend to assume that the user knows what they want to create in advance — i.e. an end-product with an initial goal. They require an initial phase in which the user must define content and behavior up-front. This could limit open-ended exploration during the creative process, when the user does not necessary have an end-goal in mind. Further, the majority of such interfaces use language input to generate or spawn content without the user in the loop. An alternative is to enable greater interactive control and definition of the behavior of the content. Our approach emphasizes flexibility and user-control during the creative process. The user can define and iterate on content at any point. Our prototype specifically supports definition of behaviors within an interactive simulation. Within this prototype, we explore language input, combined with direct manipulation, as a way of empowering the user to program and control scenes interactively.
@@ 2.2. Dynamic Sketching Interfaces
HCI researchers have extensively explored sketching interfaces for dynamic and interactive visualizations ever since the first graphical user interface (GUI) SketchPad (Sutherland, 1963) and William Sutherland’s thesis, the forerunner of the visual programming language (Sutherland, 1966). Many works use direct manipulation and sketching techniques to help users craft interactive behaviors and toolsets for illustrated animation, UI, and visual-oriented programs. For example, works by Kazi et al. (2014b; 2014a), Landay et al. (1996), Saquib et al. (2021), and Jacobs et al. (Jacobs et al., 2018) focused on mixing illustration, programming, and prototyping. Programming by-demonstration is featured in works such as K-Sketch (Davis et al., 2008) and Rapido (Leiva et al., 2021). Scratch (Resnick et al., 2009) is a well-known visual programming environment mixing game-like interactions with user-provided content and images for a playful experience. (Subramonyam et al., 2020) supports the user in forming connections between texts and concepts to learn via active diagramming. Our interface is complementarily framed around correspondences between sketches and language elements, but we use sketch-language mapping as a control mechanism, enabling the user to create interactive simulations and behaviors for open-ended exploration.
@@ 2.3. Programming-Like Interfaces
The customizability and flexibility of such a more general interface implies a need for programmability. Examples of these include real-time world simulation systems and programmable environments such as the SmallTalk programming language(Goldberg and Robson, 1983), Scratch (Resnick et al., 2009; Maloney et al., 2010) (and similar AI voice-enabled explorations like StoryCoder (Dietz et al., 2023)), Improv (Perlin and Goldberg, 1996), ChalkTalk(Perlin et al., 2018), and creative world-building games like the Little Big Planet series (Little Big Planet, 2008; Little Big Planet 2, 2011; Ross et al., 2012) and Dreams (Dreams, 2020). These encourage interactive building of scenes, games and stories. They combine with elements of interactive visual programming and drawing/sculpting with many types of content (2D, 2.5D, 3D, images). But all use explicit interfaces for programming or programming-like functionality (nodes, wires, text). We depart from explicit UI for sketching+programming-like capability, and largely replace many programming-like functionality with the use of language. Our direction explores the use of verbal, descriptive story narration together with other input modalities (i.e., touch and pen input) to create animated and interactive graphics through sketching.
@@ 3.2.3. Drawn versus non-drawn spoken content:
Although narration was used to explain the content on-screen, not all content (such as additional animations in the case of produced-content or additional details) and not all narration mapped to each other. In other words, there is content that the user doesn’t necessarily describe via language and some content might be considered unnecessary to represent. Hence, both modalities are tightly intertwined to represent the complete picture.
@@ 3.2.5. Content modification:
For the exercises, a participant would usually multitask and refine their sketches over time after initially mentioning the entity. For example, P1init added colors to their gray bird and pond sketches as they described their experience at the scene. Additionally, the participant might make verbal corrections ( e.g. P1init initially called a ”pond” a ”lake,” but self-corrected verbally). This means that speech errors are natural in real-life and we ought to provide ways to correct it (Suhm et al., 2001). As users’ intent change, they should be able to update their content.
@@ 3.2.6. Hierarchical object model:
Overall, based on our formative observations, we needed to prioritize user control over a variety of possible creative workflows. Rather than a fully-automatic solution, we needed a synthesis of user-directed input and system-feedback in support of spontaneous creative processes.
@@ 3.3. Design Goals
In brief, we envision an interface for creative exploration with interactive capability that (a) does not impose many assumptions about the user’s intent or content, (b) focuses on the process, not just on the artifact, and (c) does not require programming knowledge. Above all, the user should have control. To that end, we wanted an interface that:
D1 has minimal system assumptions
where the user controls the creative process, and the representation and behavior for sketches
@@
in that it supports improvisation, quick changes, and rapid iteration of ideas with a spontaneous feel, where operations are easily accessible and doable in any order
@@
without the need for a coding interface
@@ 4. DrawTalking
Figure 2. Interface Overview: Shown is an interface screen (from P4 in section 5). The toolbar (left) enables edit operations, e.g. copy, delete, attach/detach, save sketch. The transcript view displays the user’s speech input in an interactive panel, and the semantic diagram displays the machine’s understanding of the input. ”Speech controls” stage/confirm an action, discard input, and toggle speech recognition. ”Transcript controls” offer quick transcript text selection. The status bar displays the current color, a compass pointing ”up,” and has the pen/eraser state-change buttons and indicators. The scene shown is just before the user confirms a command for the utterance ”The character jumps on the platforms”; it selects the sketch labeled ”character” and all sketches labeled ”platform.” Tapping the ”language action” button (top-right) will stage the command and display the semantic diagram (top-left) representing the machine’s understanding of the input with selected objects displayed under their respective words; the user confirms the command by tapping again. This causes the character to jump on all of the platforms. For details on the workflow and commands, see subsection 4.2.
@@
[A mock-up of sketch characters and props playing fetch, and the user interacting by moving the characters around]A mock-up of sketch characters and props playing fetch, and the user interacting by moving the characters around. Panels 1-2 show the boy repeatedly throwing a ball into a pool of water, with the dog giving the ball to the boy afterwards, in an infinite cycle. Below the command reads, ”Over and over the boy throws the ball into the water and then the dog gives the ball to the boy.” Panels 3-4 show the user moving the boy sketch around as the command continues, forcing the dog to change its trajectory as the targets change during the simulation. It’s a demonstration of user-in-the-loop interaction.

### v4

@@ Abstract.
We introduce DrawTalking, an approach to building and controlling interactive worlds by sketching and speaking while telling stories. It emphasizes user control and flexibility, and gives programming-like capability without requiring code.
An early open-ended study with our prototype shows that the mechanics resonate and are applicable to many creative-exploratory use cases, with the potential to inspire and inform research in future natural interfaces for creative exploration and authoring.
@@
journalyear: 2024copyright: rightsretainedconference: The 37th Annual ACM Symposium on User Interface Software and Technology; October 13–16, 2024; Pittsburgh, PA, USAbooktitle: The 37th Annual ACM Symposium on User Interface Software and Technology (UIST ’24), October 13–16, 2024, Pittsburgh, PA, USAdoi: 10.1145/3654777.3676334isbn: 979-8-4007-0628-8/24/10ccs: Human-centered computingccs: Human-centered computing Graphical user interfacesccs: Human-centered computing Natural language interfacesccs: Human-centered computing Interaction techniquesccs: Human-centered computing Interactive systems and tools
@@
[This figure showcases the DrawTalking sketching + speaking controls.]This figure showcases the DrawTalking sketching + speaking controls. (1) On the left is a real-world photo of the user labeling a sketch on the digital canvas. The canvas contains a few user-drawn rectangle objects. The user narrates ”and I’m also going to have a rectangle,” which shows-up in a text transcript. Simultaneously, the user holds one of the rectangles with their finger and tapping the word ”rectangle” in the aforementioned transcript. This labels the user’s sketch with the word ”rectangle.” This is important because now the system knows what the name of the user’s object is, meaning it can now be controlled with more speech commands. (2) On the right is a mock-up of an example scenario; 3 panels. Panel 1: 3 lily pads have been drawn, as well as a frog. A caption reads a speech command: ”Forever this frog hops to a lily.” The user’s finger points to the frog to indicate which frog is being commanded. Dotted arrows indicate that the frog will now hop between the lilies once the command stops. Panel 2: Now a butterfly and some flowers have been drawn. The new command caption reads, ”The butterfly follows the frog.” Panel 3: The command caption reads ”This butterfly is slow” and the user’s finger is touching the butterfly, meaning the butterfly’s speed will now be decreased. Dotted arrows are now indicating that the butterfly has been moving around, following the frog as previously commanded. Slowing the butterfly does not require redoing the previous command.
@@ 1. Introduction
Sketching while speaking aids innovation, thinking, and communication — with applications in animation, game design, education, engineering, rapid prototyping, storytelling, and many other creative and spontaneous activities (Fan et al., 2023; Tversky, 2011). The combination enables us to think about and share anything through make-believe – including things that do not or cannot exist. We achieve this by assigning representations (sketches) to semantic concepts (objects, behaviors, properties) (Turner, 2016). For example we might suspend our disbelief so a square represents anything such as a house, a playing card, a map, a dog, or a person.
Furthermore, we often engage with an audience or conversational partner while drawing and talking, whether it’s through storytelling to one’s child or explaining concepts during a whiteboard lecture. These types of drawing and talking interactions have a spontaneous, playful, exploratory, and improvisational feel with fluid interactions. Storytelling and drawing while interacting with a standard GUI is arguably challenging if we wish to achieve this feel and fluidity. The goal, then, is to explore drawing and talking for these scenarios as an alternative, additional input scheme.
Specifically, we explore how we can take advantage of the presence of narrative to support richer human-computer interactions in creative world-building, where speech can both be used as verbal storytelling addressed to people and as a simultaneous input to the machine. We are motivated to design interactions built around this presence of narrative to enable richer human-computer interactions in creative world-building.
@@
This work is an attempt to realize a style of spontaneous interaction that seamlessly integrates sketching and talking-out-loud to build interactive, explorable worlds, thus greatly increasing our range of computational expression (Victor, 2014). In the design of these interactions, we explore the use of speech and direct input in the context of storytelling. Prior works in interactive sketching (Sutherland, 1963; Kazi et al., 2014a; Suzuki et al., 2020; Perlin et al., 2018; Saquib et al., 2021), language/AI-mediated interfaces (Winograd, 1972; Bolt, 1980; Liu et al., 2023; Xia et al., 2023; Huang et al., 2020; Xia, 2020; Subramonyam et al., 2018; Laput et al., 2013), and visual programming-adjacent interfaces or games (Sutherland, 1966; Resnick et al., 2009; Landay, 1996; Perlin and Goldberg, 1996; cycling74, 2018; Little Big Planet 2, 2011; Dreams, 2020) have laid valuable groundwork. However, they often require content to be pre-built or assume that users have predetermined goals, limiting the user’s potential to be spontaneous. Tools focus on generating a specific output rather than facilitating an ongoing creative process (Compton and Mateas, 2015). They might enforce given representations of objects (e.g. realistic or specific sketch recognition). They might feature complex UI, and in the case of programming-oriented tools, require explicit programming knowledge (whether via text, nodes, or blocks).
@@
In our prototype system, DrawTalking, users speak while freehand-sketching to create, control, and iterate on interactive visual mechanisms, simulations, and animations, with the freedom to tell stories using the same speech input as part of the narrative [^1] Informally, we distinguish between storytelling and narration: storytelling is a process achieved by mixing possibly many modalities (narration, speech, text, visuals). Narration refers to the use of speech or text — possibly, but not always for storytelling.. Via speech and direct manipulation, the user names their sketches to provide semantic information to the system, and narrates desired behaviors and rules to influence the interactive simulation, as if by explaining to others or telling a story. This is inspired by the way in which people might speak and/or write text to describe objects as they are being drawn.
The core aspect of the user-machine interaction design is the simultaneous use of speech as narrative, and as input to the system, to enable fluid storytelling and building of interactive worlds. The user narrates, points, and draws as if telling the story of their drawings to an audience. The user’s narration communicates intent to the machine by labeling (naming) or commanding (directing) objects. As a result, the user gains fluid programming-like capability through their narration, as opposed to needing to narrate while multitasking with a complex GUI and the potential audience. This allows for a natural extension to the GUI in the context of storytelling.
The second key insight is that narration and pointing as in storytelling can provide the machine with names and properties belonging to sketches, so in this storytelling context, the machine does not need to guess the user’s intended meaning behind their drawings. As a result, users can build and control interactive worlds by simple combinations of sketching and speaking while telling a story.
DrawTalking, in addition, requires no preparation or staging step, and supports the user in making changes to content and behavior anytime. By design, we balance AI-automation with user-direction such that the user is the one who chooses how to represent content, when to do operations, and what logic to define.
@@
  • DrawTalking: a novel sketching+speaking interaction with a design supporting fluid worldbuilding within the context of storytelling. It aims to balance user-agency with machine automation.
@@
  • An early qualitative study of DrawTalking that reveals its use for emergent creativity and playful ideation.
@@ 2.1. Natural Language-Adjacent Interfaces
Systems such as SHRDLU (Winograd, 1972) and Put That There (Bolt, 1980) pioneered the vision of employing natural language to communicate with computers. Due to recent advances in speech recognition and natural language understanding, the popularity of this interaction modality has exploded, and has been used in a wide range of domains. For example, VoiceCut (Kim et al., 2019) and PixelTone (Laput et al., 2013) allow users to speak short phrases or sentences to perform desired operations in image editing applications, but these applications are heavily domain-specific. Tools like WordsEye (Coyne and Sproat, 2001), Scones (Huang et al., 2020), and CrossPower (Xia, 2020) enable scene generation or content editing via language, and interfaces such as Visual Captions (Liu et al., 2023), RealityTalk (Liao et al., 2022), and CrossTalk (Xia et al., 2023) use language structure to make content appear during talks or conversations. However, many of these interfaces tend to assume that the user knows what they want to create in advance — i.e. an end-product with an initial goal. They require an initial phase in which the user must define content and behavior up-front. This could limit open-ended exploration during the creative process, when the user does not necessarily have an end-goal in mind. Further, the majority of such interfaces use language input to generate or spawn content without the user in the loop. An alternative is to enable greater interactive control and definition of the behavior of the content. Our approach emphasizes flexibility and user-control during the creative process. The user can define and iterate on content at any point. Our prototype specifically supports definition of behaviors within an interactive simulation. Within this prototype, we explore language input, combined with direct manipulation, as a way of empowering the user to program and control scenes interactively. Notably, although speech+pointing commands aren’t new (e.g. ”Put That There” (Bolt, 1980)), we focus on the use of speech input as leveraging a narrative that is required even when using standard input.
@@ 2.2. Dynamic Sketching Interfaces
HCI researchers have extensively explored sketching interfaces for dynamic and interactive visualizations ever since the first graphical user interface (GUI) SketchPad (Sutherland, 1963) and William Sutherland’s thesis, the forerunner of the visual programming language (Sutherland, 1966). Many works use direct manipulation and sketching techniques to help users craft interactive behaviors and toolsets for illustrated animation, UI, and visual-oriented programs. For example, works by Kazi et al. (2014b; 2014a), Landay et al. (1996), Saquib et al. (2021), and Jacobs et al. (Jacobs et al., 2018) focused on mixing illustration, programming, and prototyping. Programming by-demonstration is featured in works such as K-Sketch (Davis et al., 2008) and Rapido (Leiva et al., 2021). Scratch (Resnick et al., 2009) is a well-known visual programming environment mixing game-like interactions with user-provided content and images for a playful experience. texSketch (Subramonyam et al., 2020) supports the user in forming connections between texts and concepts to learn via active diagramming. Our interface is framed in a complementary way around correspondences between sketches and language elements, but we use sketch-language mapping as a control mechanism, enabling the user to create interactive simulations and behaviors for open-ended exploration.
@@ 2.3. Programming-Like Interfaces
The customizability and flexibility of such a more general interface implies a need for programmability. Examples of these include real-time world simulation systems and programmable environments such as the SmallTalk programming language(Goldberg and Robson, 1983), Scratch (Resnick et al., 2009; Maloney et al., 2010) , Improv (Perlin and Goldberg, 1996), ChalkTalk(Perlin et al., 2018), and creative world-building games like the Little Big Planet series (Little Big Planet, 2008; Little Big Planet 2, 2011; Ross et al., 2012) and Dreams (Dreams, 2020). These encourage interactive building of scenes, games and stories. They combine elements of interactive visual programming and drawing/sculpting with many types of content (2D, 2.5D, 3D, images). However, all use explicit interfaces for programming or programming-like functionality (nodes, wires, text). For example, voice-enabled AI interfaces like StoryCoder (Dietz et al., 2021) integrate a collaborative agent with which to communicate to create story sequences, but still use the block-based programming GUI as in e.g. Scratch. PUMICE (Li et al., 2019) is an AI agent specializing in defining new tasks on existing mobile GUIs via a conversation with the user. To learn unknown functionality, the agent repeatedly asks the user to demonstrate by direct manipulation with the app or to explain it by speech. Although comparable to our approach in terms of using speech+direct manipulation as a form of programming, PUMICE requires interaction with a chat-bot-like interface by design, and requires existing target applications (a different, specialized use case as opposed to ours). In contrast, we wanted to support simultaneous narration and (potentially) interaction with an audience while building a scene, with no pre-existing application or goal in-place. This leads to different interaction design criteria: we choose not to use an AI assistant because this could divert attention from an audience. Rather, the user speaks in narrative form as in storytelling and mixes direct manipulation (touch, drawing). The same input functions as both storytelling to an audience and commands to the machine, so the user does not need to address an assistant.
Overall, we depart from explicit UI for sketching+programming-like capability, and largely replace much programming-like functionality with the use of language. Our direction explores the use of verbal, descriptive story narration together with other input modalities (i.e., touch and pen input) to create animated and interactive graphics through sketching while potentially communicating to an audience as one might during teaching or storytelling.
@@ 3.2.3. Drawn versus non-drawn spoken content:
Although narration was used to explain the content on-screen, not all content (such as additional animations in the case of produced-content or additional details) and not all narration mapped to each other. In other words, there is content that the user doesn’t necessarily describe via language and some content might be considered unnecessary to represent. Hence, the modalities complement each other to represent the complete story.
@@ 3.2.4. Speed and flow:
Additionally, participants would sometimes pause to locate nested UI components (e.g. opening a color-picking menu) or switch between pen and eraser modes. To complement the GUI, speech could allow users access functionality by referring to it without needing to find it. This could help users keep the flow of their narration.
@@ 3.2.5. Content modification:
For the exercises, a participant would usually multitask and refine their sketches over time after initially mentioning the entity. For example, P1init added colors to their gray bird and pond sketches as they described their experience at the scene. Additionally, the participant might make verbal corrections (e.g. P1init initially called a ”pond” a ”lake,” but self-corrected verbally). This means that speech errors are natural in real-life and we ought to provide ways to correct it (Suhm et al., 2001). As users’ intentions change, users should be able to update their content.
@@ 3.2.6. Hierarchical object model:
Based on our formative observations, we needed to prioritize user control over a variety of possible creative workflows. Rather than a fully-automatic solution, we needed a synthesis of user-directed input and system-feedback in support of spontaneous creative processes.
@@ 3.3. Design Goals
We envision an interface for creative exploration with interactive capability that (a) is controlled via drawing and talking in the context of narration, (b) does not impose many assumptions about the user’s intent or content, (c) focuses on the process, not just on the artifact, and (d) does not require programming knowledge. Above all, the user should have control.
To that end, we wanted an interface that:
D1 makes minimal system assumptions
where the user controls the creative process, and the representation and behavior of sketches.
@@
in that it supports improvisation, quick changes, and rapid iteration of ideas with a spontaneous feel, where operations are easily accessible and doable in any order.
@@
without the need for a coding interface.
@@ 4. DrawTalking
Figure 2. Interface Overview: An interface screenshot (from P4 in section 6). The toolbar (left) enables edit operations, e.g. copy, delete, attach/detach, save sketch. The transcript view displays the user’s speech input in an interactive panel, and the semantic diagram displays the machine’s understanding of the input. ”Speech controls” stage/confirm an action, discard input, and toggle speech recognition. ”Transcript controls” offer quick transcript text selection. The status bar displays the current color, a compass pointing ”up,” and has the pen/eraser state-change buttons and indicators. The scene shown is just before the user confirms a command for the utterance ”The character jumps on the platforms”; it selects the sketch labeled ”character” and all sketches labeled ”platform.” Tapping the ”language action” button (top-right) will stage the command and display the semantic diagram (top-left) representing the machine’s understanding of the input with selected objects displayed under their respective words; the user confirms the command by tapping again. This causes the character to jump on all of the platforms. For details on the workflow and commands, see subsection 4.2 and subsection 4.4.
[…]
[A screen-shot of the DrawTalking interface, showing full canvas interface elements and semantics diagram mechanics.]A screen-shot of the DrawTalking interface, showing full canvas interface elements and semantics diagram mechanics. On bottom is the canvas containing a scene of hand-drawn sketches: left-to-right, a cartoon star character labeled ”character” and ”playable”, the surface of the moon ”floor,” below the entire scene, a rocky cliff labeled ”boulder”, spikes labeled ”spikes”, another rocky cliff called ”boulder,” two moving platforms labeled ”platform,” one above the spikes, the other after the second boulder, a flag labeled ”flag.” Around the frame of the interface, bottom-left to top-right: bottom-left: a status bar containing a horizontal color slider, and indicators for the pen control states; left: a vertical toolbar of touchable buttons, to-to-bottom (the relevant ones are): ”copy,” ”copy attached,” ”delete,” ”delete all,” ”scale,” ”flip left/right,” ”toggle show attached,” ”toggle labels,” ”pause,” ”save thing;” top-left: a rectangular region containing the semantics diagram, which looks like the simple building blocks for a sentence with objects from the scene corresponding to the nouns as small proxy objects underneath the noun blocks. The diagram reads ”character, jump, on, platform” with the character underneath the character block, and both platforms underneath ”platform” because the user has specified ”The character jumps on the platforms.” in their speech command, shown in the transcript on the right. Between the diagram and the transcript are three control buttons for manipulating the text in the transcript: ”select all,” ”selection clear,” ”selection invert,” ”label things” (which will label unlabeled objects connected to the semantics diagram with the corresponding nouns in the diagram). On the far top-right are controls for speech commands: (the relevant ones are) ”language action,” ”discard,” ”speech ignore/enable.” Top-right corner: a circular indicator colored orange when the user is speaking, grey otherwise.
@@
[An example of creating objects, naming them, commanding them, and interacting with them by moving the characters around, for a simulation of a game of fetch between a boy and his dog.]An example of creating objects, naming them, commanding them, and interacting with them by moving the characters around, for a simulation of a game of fetch between a boy and his dog. Two sub-panels a) and b). a): the process of creating and naming the necessary objects. Panel 1: a ball has been drawn and the user is in the process of drawing a boy with the pen, the boy has the name ”boy” underneath. Caption: ”This is a boy”; Panel 2: The boy is fully drawn and has the name ”boy” below. The ball is shown tapped by the user’s finger and now has the name ”ball” underneath it; caption ”and this is a ball; Panel 3: a drawing of a pool of water is added, the user’s hand is holding the water, caption: ”and I’m going to draw some water too, the user’s pen is holding the word water, creating an arrow between the finger-held water and the word ”water.” A blue link points from ”water” to the water sketch, establishing the link, and now naming the water sketch with the text ”water” below the sketch; Panel 4: now the touch+pen is done between a new ”dog” sketch and the word ”dog” in the remainder of the user input caption: ”Lastly I’m going to draw myself a dog.” b): Panels 1-2 show the boy repeatedly throwing a ball into a pool of water, with the dog giving the ball to the boy afterwards, in an infinite cycle. Below the command reads, ”Over and over the boy throws the ball into the water and then the dog gives the ball to the boy.” Panels 3-5 show the user moving the boy sketch around as the command continues, forcing the dog to change its trajectory as the targets change during the simulation. The full example is a demonstration of user-in-the-loop interaction.